4. +-1 points TanApMath7 2.5.022. My Notes Ask Your Teache Find the break-even point for the firm whose cost function C and revenue function R are given. C(x)-13x+7,000; R(x)-20x (x,y)-1000, 5. + -1 points TanApMath7 2.5.025. My Notes Ask Your Teache An office building worth $2 million when completed in 1997 was depreciated linearly over 50 years; that is, the book value of the building decreases at a constant rate, so that at the end of 50 years the book value is zero. What was the book value, in dollars, of the building in 20131? What will be the book value, in dollars, in 2020? What will be the book value, in dollars, in 2022? Your work in question(s) 4 will also be submited or saved. Submit Assignment Save Assignment Progress
Answer #1


4)
For break even:
C(x) = R(x)
13x + 7000 = 20x
7x = 7000
x = 1000

at x = 1000,
R(x) = 20x = 20000 = C(x)

Answer: (1000, 20000)

5)
Let x be number of years after 1997
Let y be the value in dollars
(x1,y1) = (0,2000000)
(x2,y2) = (50,0)
So,
slope = (y2-y1)/(x2-x1)
= (0, 2000000)/(50-0)
= -40,000
Equation:
(y-y1) = slope (x-x1)
(y-2000000) = -40000 (x-0)
y = -40,000x + 2000000

a)
Here x = 2013 - 1997 = 16
y = -40,000x + 2000000
= -40,000*16 + 2000000
= 1,360,000
Answer: 1,360,000

b)
Here x = 2020 - 1997 = 23
y = -40,000x + 2000000
= -40,000*23 + 2000000
= 1,080,000
Answer: 1,080,000

c)
Here x = 2022 - 1997 = 25
y = -40,000x + 2000000
= -40,000*25 + 2000000
= 1,000,000
Answer: 1,000,000